User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!


Hobart 26m

Module: HOB+1007/16000, start pos 7401.6 GB

  • 1730 UT - Experiment started on-time. Formatter/maser delay offset larger than normal (~2.8 us), it seems stable (+-0.05 us) with clkoff values stable. As the proc file does not call maserdelay I was unsure if this is normal. Called Warren and he is out at the telescope having a look if things are working correctly. At this stage it is being left as is. (Tiege)
  • 2122 UT - Schedule halted to put correct receiver on axis. Prior to this, the 22 GHz receiver was installed and no valid data would have been recorded. (JMc)
  • 2136 UT - First useful data - previous scan not reached in time after Rx change, (JMc)
  • 0606 UT - Pressure and temperatures too high - alarms going off. (25 microns Hg, 20K stage at 42K, and 70K stage at 85K) called Mt. Pleasant and Jim is restarting the vacuum pump. (Bryn)
  • 0643 UT - Temperatures still high in 20K and 70K stages (41K and 78.8K), but are slowly falling. (Bryn)
  • 0747 UT - 70K stage back at 65K. 20K stage at 35K (Bryn)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/rd1705.txt · Last modified: 2017/04/20 07:48 by Elizabeth Iles